### v1

# Assessing Logical Reasoning Capabilities of Encoder-Only Transformer Models

Abstract
Logical reasoning is central to complex human activities, such as thinking, debating, and planning; it is also a central component of many AI systems as well. In this paper, we investigate the extent to which encoder-only transformer language models (LMs) can reason according to logical rules. We ask whether those LMs can deduce theorems in propositional calculus and first-order logic; if their relative success in these problems reflects general logical capabilities; and which layers contribute the most to the task. First, we show for several encoder-only LMs that they can be trained, to a reasonable degree, to determine logical validity on various datasets. Next, by cross-probing fine-tuned models on these datasets, we show that LMs have difficulty in transferring their putative logical reasoning ability, which suggests that they may have learned dataset-specific features, instead of a general capability. Finally, we conduct a layerwise probing experiment, which shows that the hypothesis classification task is mostly solved through higher layers.
1 Introduction
Transformer models are remarkably effective at a wide range of natural language processing (NLP) tasks, such as question answering, summarization, and text generation. By and large, these abilities are the result of specific training processes, where a language model is fine-tuned on a task-specific dataset. Curiously, encoder-only transformer models (LMs) also exhibit implicit linguistic and cognitive abilities for which they were not directly supervised. Such LMs have been shown to encode information on tense and number (Conneau et al., 2018), anaphora and determiner-noun agreement (Warstadt et al., 2020), semantic roles (Tenney et al., 2019b), syntactic dependencies (Manning et al., 2020), relational knowledge (Dai et al., 2021), and spatiotemporal representation (Gurnee and Tegmark, 2023).
Hence, it is worth investigating whether encoder-only transformer models demonstrate logical reasoning capabilities as well. Given that logical reasoning plays a fundamental role in intelligence, human or otherwise (Russell and Norvig, 2010), it is essential to explore if these models exhibit such capabilities. Concerning NLP, understanding the structure of LMs’ logical reasoning may enable us to discern the extent to which their inferences and arguments arise from reasoning rather than purely associative memory. This understanding is crucial for developing mechanisms that facilitate the generation of factual text from non-structured databases.
The goal of this paper is, thus, to assess the ability of encoder-only transformer models to reason according to the rules of logic — understood here as deductive arguments expressable in propositional calculus or first-order logic. We examine three main questions throughout this paper: i) Can enconder-only transformer models perform logical reasoning tasks?; ii) How general is this ability?; and iii) What layers better contribute to solving these tasks?
Section 2 reviews the work on transformers’ logical reasoning abilities, as well as the function of probing in uncovering latent knowledge. Next, we gather and describe four datasets grounded on logical deduction (sec. 3). In a first batch of experiments, we conduct a systematic comparison of encoder-only transformer models on these datasets (sec. 4). Section 5 then investigates whether the performance in the previous task could be attributed to some general ability and whether the reasoning learned from one dataset could be transferred to a similar dataset. Finally, we perform a layer-wise probing to understand which layers are responsible for solving logical deduction problems (sec. 6). [^1] Data, code, and complete results are available at https://github.com/paulopirozelli/logicalreasoning.
2 Related Work
Logical Reasoning in Transformer Models
Transformer models are powerful enough to solve logical reasoning tasks expressed in natural language (Clark et al., 2020; Zhang et al., 2022; Tian et al., 2021; Han et al., 2022). Yet, it is not clear if these models have actually mastered logical reasoning. LMs seem to inevitably rely on statistical artifacts to deduct theorems, rather than in general, rule-based relationships (or something akin to that) (Zhang et al., 2022). They use shortcuts to solve hypothesis classification problems, leading to vulnerabilities in reasoning (e.g., LMs are fooled when hypotheses appear within rules), and making them susceptible to irrelevant (logically-consistent) perturbations (Gaskell et al., 2022).
Studies focused on functional words close to logical operators have identified similar shortcomings in LMs’ reasoning capabilities. Transformers struggle to deal with negation, predicting similar probabilities to a sentence and its negation (Kassner and Schütze, 2020; Hossain et al., 2020). Kalouli et al. (2022) extend these findings to conjunction, disjunction, and existential and universal quantification, showing that expressions associated with these operations are frequently dominated by semantically rich words. Transformers also fail in modeling semi-functional-semi-content words in general, such as quantifiers and modals (Sevastjanova et al., 2021). Tangentially to logic, Dankers et al. (2022) show that transformer models do not always properly handle compositionality: sometimes a translation is more local than desired (treating idioms as regular constructions), sometimes it is excessively global (paying attention to irrelevant parts of the sentence). Moreover, there is evidence that the way LMs compose sentences does not align well with human judgment (Liu and Neubig, 2022).
Large LMs have also been assessed for their logical reasoning capabilities. Despite their impressive achievements in numerous tasks, these models still struggle with multi-step and compositional problems (Rae et al., 2021; Dziri et al., 2023). Although good at individual deductions, large LMs struggle with proof planing: when many valid proof steps are available, they often take the wrong path (Saparov and He, 2023). Large LMs also appear to suffer from human-like biases in logical tasks: they perform significantly worse when the semantic content is too abstract or conflict with prior knowledge (Dasgupta et al., 2022; Tang et al., 2023).
Probing Tasks
It is usually hard to tell which data features are employed by an LM in solving a problem, and whether the LM is organizing its operations in human-like ways. By definition, fine-tuning is not a viable option to get insights about this process, as it requires reparametrizing the network; as a consequence, it is not possible to distinguish what has been learned in the task-oriented training and what was there from the beginning.
Probing tasks aim at recovering latent knowledge from the parameters of an LM: a dataset that encodes a particular property (e.g., POS) is used to train a classifier (the probe), taking the representations produced by the LM as inputs for the probe (Belinkov, 2022). As the LM is not further trained on the task, the performance of the probe depends on the information on that property been previously encoded by the model. Probing may be used to examine various components of LMs, such as embeddings, attentions or feedforward computations.
Transformer models have been extensively studied through probing (Belinkov, 2022). Most attention has been given to BERT, which gave rise to a large literature on the properties encoded by this model (Rogers et al., 2020). RoBERTa has also been studied in some detail through probing; e.g., what abilities that model learns during training (Zhang et al., 2021; Liu et al., 2021) and its knowledge of semantic structural information (Wu et al., 2021).
Figure 1: Examples of logical reasoning arguments. The argument on the top is from FOLIO, a manually-written dataset; the one in the bottom is from RuleTaker, a dataset that uses a semi-synthetic approach.
3 Logical Reasoning Datasets
Table 1: Main features of the logical reasoning datasets. FOL stands for first-order logic, PC for propositional calculus, and CI for conjunctive implication. Labels are: FOLIO (False, True, Unknown), LogicNLI (Contradiction, Entailment, Neutral, Paradox), RuleTaker (False, True), SimpleLogic (False, True). The average number of premises and the average number of words per argument is calculated for the training set. Appendix B shows the full label distribution.
Dataset Size (train/val/test) Scope Type Label Ave. Premises Ave. Words
FOLIO 1003 / - / 204 FOL Manual 3 5.23 64.01
LogicNLI 16000 / 2000 / 2000 FOL Semi-synt. 4 24 211.86
RuleTaker 27363 / 3899 / 7793 FOL (CI) Semi-synt. 2 16.30 100.06
SimpleLogic 11341 / 1418 / 1417 PC (CI) Synthetic 2 60.76 467.45
In assessing logical reasoning, we restricted our analysis to datasets related to propositional calculus (PC) and first-order logic (FOL); FOL being an extension of PC that includes predicates and quantifiers. These logical systems offer a powerful formalism that balances simplicity and expressivity in representing and reasoning about statements and relationships. They are suitable to capture a wide range of knowledge and to formalize a large number of domains.
In addition to being expressable in PC or FOL, we selected datasets that satisfied three other properties: i) observations had to be self-contained and to require no prior knowledge; ii) sentences needed to have correspondent translations in both logical formalism and natural language; and iii) hypotheses had to be declarative sentences. The first property lets one decouple logic from background knowledge, in order to assess pure reasoning capabilities; this prevented the LMs’ from using their implicit knowledge base. For that reason, we only considered pure logical datasets. Thus, we did not take into account other forms of reasoning, such as scientific reasoning (e.g., Clark et al. 2018; Lai et al. 2017; Yu et al. 2020); mathematical reasoning (e.g., Cobbe et al. 2021; Patel et al. 2021; Hendrycks et al. 2021); counterfactual reasoning (e.g., Qin et al. 2019; Yang et al. 2020a; O’Neill et al. 2021); planning (e.g., Valmeekam et al. 2023); inductive reasoning (e.g., Sinha et al. 2019; Tang et al. 2023); and abductive reasoning (e.g., Tang et al. 2023). The second property, the translation in a logical formalism, is what allows one to determine unambiguously the logical relationship between premises and hypotheses, while the natural language counterpart allows us to probe the LMs. Hence, we excluded datasets that lacked natural language translations, such as LTL (Hahn et al., 2020). The last property excluded QA datasets as well (e.g., Liu et al. 2020; Talmor et al. 2019), as they required the understanding of several types of questions (e.g., why, what) that are entangled with semantic and contextual knowledge (e.g., that “Alice” is the name of a person).
In the end, four logical reasoning datasets were selected: FOLIO (Han et al., 2022), LogicNLI (Tian et al., 2021), RuleTaker (Clark et al., 2020), and SimpleLogic (Zhang et al., 2022); some examples of arguments can be seen in Figure 1. These datasets cover a wide range of variations in terms of construction (manual, semi-synthetic, synthetic), alignment with common sense, linguistic variability, and scope (PC, full or partial FOL). For instance, FOLIO is human-written; LogicNLI and RuleTaker both use a template which is then manually edited; and SimpleLogic is fully synthetic. FOLIO and LogicNLI encompass the full spectrum of FOL; RuleTaker is expressable in FOL, but only covers negation and conjunctive implications; [^2] Conjunctive implications are arguments of the form ($fact.[\land fact.]^{*}rule.[\land rule.]^{*}\Rightarrow hypothesis$). and SimpleLogic is restricted to conjunctive implications in PC. As regarding the number of labels, RuleTaker and SimpleLogic have True and False labels; FOLIO includes an Unknown label; and LogicNLI admits a fourth possibility, Paradox, where both a hypothesis and its negation can be deduced from the premises.
[…]
4 Testing LMs for Hypothesis Classification
The performance of LMs in logical reasoning, and in particular in the datasets described in the previous section, has been studied in scattered experiments, without a clear unified context that allows direct comparison. Due to this, we fine-tuned a wide range of pretrained encoder-only LMs in those datasets, applying them to a hypothesis classification task; i.e., determining the logical relationship between a set of premises and a conclusion. These results represent a small yet valuable contribution to the existing literature.
To ensure comprehensive coverage, eight families of encoder-only models were assessed: DistilBERT (Sanh et al., 2020), BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), Longformer (Beltagy et al., 2020), DeBERTa (He et al., 2020), AlBERT (Lan et al., 2019), XLM-RoBERTa (Conneau et al., 2020), and XLNet (Yang et al., 2020b). The full list of models is displayed in the left column of Table 2.
Table 2: Accurary comparison among several encoder-only Language Models (LMs) for the hypothesis classification task across various datasets: FOLIO (FL), LogicNLI (LN), RuleTaker (RT), and SimpleLogic (SL). The table depicts different model types in the left column. Results for RoBERTa-large, the LM picked for subsequent analyses, are in bold. Best and worst scores for each dataset are highlighted in blue and red, respectively.
FL LN RT SL
DistilBERT 50.98 36.20 86.55 93.58
BERT-B 52.45 48.75 98.91 92.31
BERT-L 61.76 42.50 99.94 91.88
RoB-B 58.82 62.90 98.04 92.87
RoB-L 64.71 72.70 99.78 90.83
Longf-B 62.75 58.05 99.92 94.21
Longf-L 62.25 74.15 99.94 92.37
DB-XS 54.41 65.30 99.81 91.67
DB-S 53.43 59.65 95.23 93.23
DB-B 60.78 66.70 99.81 93.72
DB-L 64.71 84.70 99.97 93.72
DB-XL 62.25 25.00 99.99 49.96
DB-XXL 71.57 25.00 50.02 50.04
AlB-B 57.35 66.80 99.91 92.17
AlB-L 56.37 66.20 99.88 91.53
AlB-XL 38.73 65.10 99.97 90.05
AlB-XXL 56.86 93.90 99.94 92.24
XLM-B 55.88 45.20 97.64 91.74
XLM-L 62.75 65.45 99.95 91.74
XLNet-B 58.33 55.00 99.19 94.28
XLNet-L 58.33 71.40 99.86 92.94
Larg. class 35.29 25.00 50.02 50.03
To fine-tune our models, we used a linear layer with dimension (embedding-length, labels) and dropout of 0.5. As input for the classification head, we used the last hidden embedding of the [CLS] token, as usual for classification tasks in NLP. Sequences were padded and truncated to the maximum length allowed by each LM. Models were trained for up to 50 epochs, using early stopping with a patience of 5 epochs. We used Adam as our optimizer ($\beta$ = [0.9, 0.999] and no weight decay) and experimented with two different learning rates (1e-5 and 1e-6). Models were chosen based on the loss for the validation set, and results reported for the test set. The only exception is FOLIO, for which only the validation set is publicly available; results are thus reported for the latter. We report accuracy as the standard metric, as classes in the datasets are balanced. Table 2 displays the best result achieved in each case. The best and worst scores for each dataset are highlighted in blue and red, respectively. We also provide a largest class baseline for comparison. [^3] Complete results, including the losses for each set and the F1-score (micro, macro, and weighted) for the test set, can be found at the project’s Github page.
Results show that the LMs were able to classify hypotheses with reasonable success. All of them came close to solve RuleTaker, with all but DeBERTa-XL achieving an accuracy in SimpleLogic above 90%. Results were comparatively lower for FOLIO and LogicNLI, but still LMs considerably surpassed the largest class baselines. A weaker performance observed in these datasets was expected, given their greater language variability and broader logical scope; and in the case of FOLIO, smaller size as well. Overall, the encoder-only transformer models worked well as soft reasoners (Clark et al., 2020), being able to successfully deduce theorems from premises expressed in natural language. Noteworthy exceptions were the performances of AlBERT-XL in FOLIO; DeBERTa-XL and -XXL in LogicNLI and SimpleLogic; and DeBERTa-XXL in RuleTaker.
5 Cross-Probing Fine-Tuned LMs
LMs can successfully solve hypothesis classification tasks that involve logical reasoning. However, it raises the question: has this logical reasoning ability been acquired during the fine-tuning stage, or was it present from pretraining? Most importantly, have LMs truly developed a generalized logical reasoning capability? To examine this, we run a cross-probing task, where the LMs fine-tuned on our four logical datasets, as well as a pretrained LM, are probed on these same datasets.
Given the large number of possible tests, we restricted our investigation to a single model, RoBERTa-large, so as to dig deeper on it. This LM demonstrated a suitable balance between performance, consistency among datasets, and training time in the fine-tuning tests.
For the cross-probing task, we took the best fine-tuned RoBERTa-large model for each dataset and removed their classification heads, leaving just the transformer blocks, as in a pretrained model. We then trained a probe on the LMs’ final representations, while freezing the models’ bodies; we also probed a pretrained RoBERTa-large model. The same training policies from the fine-tuning step were followed in this stage. Two different classifiers were tested as probes:
  • 1-layer A single affine transformation is applied to the last hidden state of the first token (the [CLS] token). The classification head has shape (1024, labels). We used a dropout of 0.5 before the classifier.
  • 3-layer The [CLS] embedding passes through three consecutive layers of shape (1024, 256), (256, 64), and (256, labels), respectively. We used a dropout of 0.5 in between linear layers and ReLU as the activation function.
Both probes led to similar results. We take this as strong evidence that the knowledge used in the logical reasoning tasks can be linearly recovered from the internal representations of RoBERTa-large.
Table 3: Results for the cross-probing task: On the left, we present the best fine-tuned RoBERTa-large models for each dataset. The datasets used for training and testing the probes are listed on the top. We report only the best result for each probe. Blue cells indicate the probe of a fine-tuned model on the same dataset. In parentheses, we denote the percentage difference from the pretrained model. “Largest class” refers to the score for selecting the class with the highest frequency in the training set.
FOLIO LogicNLI RuleTaker SimpleLogic
Pretrained 32.88 25.54 50.01 61.19
FOLIO 55.05 (+67.42) 27.28 (+6.81) 60.74 (+21.45) 62.82 (+2.66)
LogicNLI 36.60 (+11.31) 67.95 (+166.05) 69.37 (+38.70) 62.06 (+1.42)
RuleTaker 40.32 (+22.62) 36.01 (+40.99) 99.44 (+98.84) 62.89 (+2.77)
SimpleLogic 32.88 (0) 25.44 (-0.39) 51.02 (+2.01) 92.35 (+50.92)
Larg. class 35.29 25.00 50.02 50.03
Table 3 displays the best results attained in the cross-probing task. The left column lists the best RoBERTa-large fine-tuned models from the previous experiment, while the remaining columns represent the datasets they were probed against. The blue cells along the diagonal indicate instances where models were probed on the same datasets they were initially fine-tuned on. Percentage differences in accuracy relative to the pretrained case are reported in parentheses.
As expected, the fine-tuned model for a specific dataset performed better in that same dataset, albeit less than in the fine-tuning scenario. This makes sense, since in fine-tuning, both the model’s body and head are optimized, whereas in a probe the head is in charge of all the learning. These results serve as a sanity check that our probing is working.
The first row in Table 3 contains the scores for the probes with pretrained RoBERTa. Accuracy levels for LogicNLI and RuleTaker closely resemble the largest class baseline, while the result for FOLIO falls below its corresponding baseline. Pretrained RoBERTa only helped to solve SimpleLogic, a dataset constrained to conjunctive implication with minimal language variation. Its pretraining scheme, dynamic masked language modeling, did not equip it with adequate logic-like knowledge to solve more complex reasoning problems without specific training.
Clearly, if logical reasoning was not in pretrained RoBERTA, we would expect it to have been acquired in the fine-tuning process. After all, RoBERTa was able to solve the hypothesis classification tasks after training. When analyzing the results of the cross-probing task, however, we may have some doubt as to whether a general logical reasoning ability in fact emerged from fine-tuning.
In general, the fine-tuned LMs showed limited transferability when probed on different datasets. Although some gain was achieved as compared to the pretrained model, they remained well below what the LM fine-tuned on the same dataset obtained. SimpleLogic presents a singular case. The LMs fine-tuned on the other datasets performed similarly to the pretrained model in this dataset. This is despite SimpleLogic covering only a subset of propositional calculus, a domain included in those datasets. Conversely, the LM originally fine-tuned on SimpleLogic did not exhibit improved performance on other datasets, indicating a lack of acquired general logical reasoning ability during its training.
Two main conclusions can be taken from this experiment:
  1. 1.
    If the difference in accuracy between pretrained RoBERTa and the largest class baseline is indicative of the amount of logical reasoning contained in this LM, then pretrained RoBERTa’s logical reasoning abilities are very limited.
  2. 2.
    If the difference in accuracy between fine-tuned LMs (when applied to different datasets) and the pretrained RoBERTa model is indicative of the amount of logical reasoning they acquired in the fine-tuning process, then these LMs have acquired little or no general logical reasoning capability — suggesting that they mostly learned statistical features of the datasets. This is in line with other findings for transformer models (Zhang et al., 2022; Dziri et al., 2023).
6 Inspecting Fine-Tuned Models Layerwise
We now address another question: which parts of the LMs are more capable of solving logical reasoning tasks? To answer that, we probe the different layers of the fine-tuned RoBERTa-large models using the same datasets they were trained on. Our goal is to identify which layers are more effective in deducing hypotheses. Similar to the previous experiment, the fine-tuned LMs were frozen, and a classifier was trained on top of each layer. However, in this case, the fine-tuned LMs were only probed using the datasets they were previously trained on. The same classifiers from the last experiment were tested. They were positioned on top of the [CLS] token of each layer, with 25 layers in total (24 transformer blocks plus the initial embedding layer). We adopted the same configurations from the previous tests: models were trained with early stopping for up to 50 epochs (patience of 5 epochs) based on the validation loss, using two learning rates (1e-5 and 1e-6).
Figure 2: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 1-layer classifier and a learning rate of 1e-6.
[…]
A similar behavior is exhibited by all models. They remain close to the pretrained baseline in the low and mid layers. Accuracy then grows rapidly in the final layers, achieving a performance equal to the cross-probing baseline.
How does that compare to other types of knowledge found on transformer models? Literature on transformers show that surface information is mostly captured by lower layer information, e.g., sentence length (Jawahar et al., 2019); middle layers are responsible for processing syntactic knowledge, e.g., syntax trees (Hewitt and Manning, 2019); and higher layers are responsible for task-specific functions (Hao et al., 2019) and contextual representations (Ethayarajh, 2019).
In our layerwise probing, higher layers were the only layers ones that were able to solve the hypothesis classification task better than a pretrained model. This suggests that the knowledge acquired in the fine-tuning training was connected to dataset-specific features, rather than general representations. It also explains why the information was not transferable among datasets.
SimpleLogic was the only case that presented a growth in the initial layers. This behavior may indicate that the dataset is solvable through the use of some heuristics based on shallow statistical features, such as the number of premises, as discussed by Zhang et al. (2022).
7 Conclusion
Logical reasoning is a valuable ability that humans use in thinking, arguing, and planing; it is also a core component of many AI systems. We investigated here the role of logical reasoning in encoder-only transformer models. By gathering a number of logical reasoning datasets, we saw that LMs can be trained to perform complex logical tasks with relative success. However, upon closer inspection, doubts arose regarding whether these models have truly learned to reason according to logical rules. First, by probing a pretrained RoBERTa-large model with logical reasoning datasets, it became apparent that this language model did not possess intrinsic logical reasoning abilities. Second, models fine-tuned on one dataset struggled to generalize well to others, even within the same domain. This observation suggests that these language models did not acquire robust logical reasoning capabilities. Third, the knowledge necessary to solve logical reasoning tasks seems to emerge primarily at higher, more contextual layers, which appear to be linked to statistical features of the datasets rather than deeper representations.
8 Limitations
We run experiments for a large variety of LMs in Section 4; however, due to space and time constraints, we focused on RoBERTa-large for our analysis in Sections 5-6. While we expect the same behavior to appear in other encode-only transformer models, further tests are needed to verify whether conclusions can be reliably extended to them. We have not explored decoder nor encoder-decoder models either, as those are better tested through behavioral tests, rather than structural probes. Furthermore, it could be the case that robust logical reasoning is an emergent ability only manifested in large language models (Wei et al., 2022).
In a different direction, other types of representations could be analyzed in relation to logical reasoning, such as attentions and feedforward computations. Further work should also concentrate on other types of logical formalism.
Appendix A Datasets
In this appendix, we describe the four logical reasoning datasets in more detail. Table 4 indicates the sources from where they were obtained.
FOLIO
(Han et al., 2022) is a human-annotated dataset for FOL reasoning problems. Logically-sound contexts were generated in two ways: in the first, annotators created contexts from scratch, based on random Wikipedia pages; in the second, a template of nested syllogisms was used, from which annotators then replaced the abstract entities and categories by nouns, phrases or clauses, as to make the text to reflect real-life scenarios. Next, the authors verified the alignment between natural language sentences and FOL formulas and added implicit commonsense knowledge as premises. After that, they verified the syntactic validity and label consistency of FOL formula annotations with a FOL prover. Finally, sentences were reviewed for grammar issues and language fluency. Only train and validation tests are available, so we used the latter for reporting tests. Hypotheses can be True, False, or Unknown.
LogicNLI
(Tian et al., 2021) is a FOL dataset created through a semi-automatic method. A set of logical templates was defined and then filled by subjects and predicates sampled from predefined sets. Next, manual edits were made to correct grammatical errors and resolve semantic ambiguities. Hypothesis are classified as Entailment, Contradiction, Neutral, and Paradox. A paradox is defined as a situation where both a sentence and its contrary can be inferred from the premises. We used the standard version of the dataset, which encompasses all labels.
RuleTaker
(Clark et al., 2020) is a logical reasoning dataset in which rules are conjunctive implications. Predicates may be negated and facts may be either attributes (which assign properties to entities) or relations (which relate two entities). We used the ParaRules version, where rules and facts were paraphrased by crowdworkers into more natural language; paraphrased constructions were then combined to form new templates. We also used the updated version of RuleTaker (problog), which eliminated some world model inconsistencies. Hypothesis can be True, when a hypothesis follows from the premises, and False otherwise (closed-world assumption, CWA).
SimpleLogic
(Zhang et al., 2022) is similar to RuleTaker, only supporting conjunctive implication (facts are simply conjunctive implications with empty antecedents). Language variance is virtually removed by using a fixed template for translating FOL into natural language and by the use of a small random list of words as predicates. Argumentative complexity is limited by setting thresholds for input length, number of predicates, and reasoning depth. We reconstructed the original template based on the examples given in the paper. For our tests, we used the RP Balanced version. We undersampled the largest class (True) to obtain the same number of observations as for the False class. Labels can be True and False (CWA).
Dataset Source
FOLIO https://github.com/Yale-LILY/FOLIO
LogicNLI https://github.com/omnilabNL
/LogicNLI
RuleTaker https://allenai.org/data/ruletaker
SimpleLogic https://github.com/joshuacnf
/paradox-learning2reason
SNLI https://nlp.stanford.edu
/projects/snli/
MultiNLI https://cims.nyu.edu
/~sbowman/multinli/
SciTail https://huggingface.co
/datasets/scitail
Table 4: Sources for the datasets.
Appendix B Label distribution
Table 5 shows the label distribution for the datasets used in the paper. The row above provides the number of observations per label, and the row below shows their relative percentage. Labels: FOLIO: False, True, Unknown. LogicNLI: Contradiction, Entailment, Neutral, Paradox. RuleTaker: False, True. SimpleLogic: False, True.
Train Validation Test
FOLIO 286 / 388 / 329 - 63 / 72 / 69
(28.51% / 38.68% / 32.80%) - (30.88% / 35.29% / 33.82%)
LogicNLI 4000 / 4000 / 4000 / 4000 500 / 500 / 500 / 500 500 / 500 / 500 / 500
(25% each) (25% each) (25% each)
RuleTaker 13666 / 13697 1946 / 1953 3895 / 3898
(49.94% / 50.05%) (49.91% / 50.09%) (49.98% / 50.02%)
SimpleLogic 5696 / 5645 683 / 735 709 / 708
(50.22% / 49.77%) (48.16% / 51.83%) (50.03% / 49.96%)
Table 5: Label distribution for logical reasoning datasets.
Appendix C Laywerwise probing
For the layerwise probing (Sec. 6), we tested two different classifiers (1-layer and 3-layer) and two learning rates (1e-6 and 1e-5). Figure 2 above displayed the results for the 1-linear classifier and 1e-6 learning rate. The figures below display the results for the other probes. Figure 3 provides the graphs for the 3-layer classifier and learning rate of 1e-6; Figure 4 provides the graphs for the 1-layer classifier and learning rate of 1e-5; and Figure 5 provides the graphs for the 3-layer classifier and learning rate of 1e-5.
Figure 3: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 3-layer classifier and a learning rate of 1e-6.
Figure 4: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 1-layer classifier and a learning rate of 1e-5.
Figure 5: RoBERTa-large models fine-tuned on FOLIO, LogicNLI, RuleTaker, and SimpleLogic, and probed for the same datasets layerwise. The pretrained baselines are indicated by the gray lines, and the values achieved in the cross-probing task are represented by the black lines. The colored bars indicate the change in accuracy from the previous layers. The probing is performed with a 3-layer classifier and a learning rate of 1e-5.